A rocket sled for testing equipment under large accelerations starts at rest and accelerates according to the expression a = (4.3 m/s 3 )t + (7.7 m/s 2 ). How far does the sled move in the time interval t = 0 to t = 2.4 s ? Answer in units of m.
Answer #1

v = \small \int a dt
v = 2.15t2 + 7.7t + vo

AT t=0, v = 0+0+vo = 0 => vo = 0
v = 2.15t2 + 7.7t

x = \small \int_{0}^{2.4} v dt
x = (2.15 / 3)t3 + (7.7 / 2)t2 (evaluate over 0 ≤ t ≤ 2.4)
x = 0.716(2.4)3 + 3.85(2.4)2
x = 32.0832 m